Circular No. 8660

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006D, 2006E, 2006F, 2006G, 2006H Two apparent supernova have been reported from unfiltered CCD ex-

posures: 2006G by M. Baek and W. Li (LOSS/KAIT; cf. IAUC 8658), and 2006H by T. Puckett and G. Sostero (0.60-m automated supernova patrol telescope; cf. IAUC 8658).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006G Jan. 13.15 1^{h}24^{m}35.^{s}83 + 1^{o}43′15′′.8 17.5 32′′.4 E, 38′′.3 S
2006H Jan. 15.05 3 26 01.49 +40 41 42.5 17.0 4′′.7 E, 2′′.2 N

Additional magnitudes of 2006G in NGC 521: 2005 Nov. 22.21 UT, [19.5; Dec. 14.20, [18.0; 2006 Jan. 13.22, 17.5; 14.12, 17.3. Additional approximate unfiltered CCD magnitudes of 2006H, whose offsets above are with respect to the center of a galaxy that itself is near UGC 2730: 2005 Sept. 23, [19.0 (Puckett); 2006 Jan. 15.77, 16.9 (G. Sostero and S. Garzia on multiple frames taken with a 0.25-m reflector). Note that the α offset for 2006H corrects that on CBET 367. Supernovae 2006D and 2006E (cf. IAUC 8658) are type-Ia events (cf. CBET 366), while 2006F (cf. IAUC 8658) is a type-Ib supernova and 2006G is a type-II supernova (cf. CBET 368).

COMET P/2000 EC_{98} = (60558)

Further to IAUC 8656, K. Kadota (Ageo, Saitama, Japan, 0.25-m reflector) reports a condensed, bright coma of diameter 1′ (total mag 14.8) and no tail on CCD images taken on Jan. 7.8 UT (communicated via S. Nakano). R. Ligustri (Talmassons, Italy, 0.35-m reflector) finds an asymmetric coma of size 1′.3 × 0′.9 extended towards the east in unfiltered CCD images taken on Jan. 9.17 UT. G. Sostero and E. Guido (remotely using a 0.25-m reflector near Mayhill, NM) report that R_{c}-band images taken on Jan. 10.49 show an uncondensed coma with diameter ∼ 0′.9 and total mag 14.1, appearing slight elongated in the north-south direction. Visual totalmagnitude estimates: Jan. 8.78, 14.4 (S. Yoshida, Ibaraki, Japan, 0.40-m reflector); 11.53, 14.4 (A. Hale, Cloudcroft, NM, 0.41-m reflector).

COMET 73P/SCHWASSMANN-WACHMANN J. A. Farrell corrects his separations and position angles as reported on

IAUC 8659: Jan. 6.48831 UT, 1573′′.7 ± 0′′.2 in p.a. 297^{o}.78 ± 0^{o}.03; 7.51252, 1598′′.9 ± 0′′.6 in p.a. 297^{o}.80 ± 0^{o}.03; 9.39304 (through high clouds), 1644′′.8 ± 1′′.3 in p.a. 297^{o}.83 ± 0^{o}.05.

2006 January 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green